# Unified Performance Control for Non-Square Nonlinear Systems with Relaxed Controllability

### v1

@@ Abstract
In this paper, we investigate the problem of unified prescribed performance tracking for a class of non-square strict-feedback nonlinear systems in the presence of actuator faults under relaxed controllability conditions. By using a skillful matrix decomposition and introducing some feasible auxiliary matrices, a more generalized controllability condition than the current state of the art is constructed, which can be applied to both square and non-square nonlinear systems subject to actuator faults and unknown yet time-varying control gain. Incorporating the relaxed controllability conditions and the uniform performance specifications into the backstepping design procedure, a prescribed performance fault-tolerant controller is developed that can achieve different performance demands without modifying the controller structure, which is more flexible and practical. In addition, the destruction of the system stability by unknown auxiliary matrices and unknown nonlinearities is circumvented by embedding the available core information of the state-dependent uncertainties into the design procedure. Both theoretical analysis and numerical simulation demonstrate the effectiveness and benefits of the proposed method.
@@
Non-square nonlinear systems, controllability condition, prescribed performance, unknown control gain, actuator fault.
@@ I Introduction
Inspired by the above discussion, in this paper, we focus on solving the problem of prescribed performance control (PPC) of strict-feedback MIMO systems under a more relaxed controllability condition. Unlike the existing literatures on PPC for MIMO systems (e.g., [8, 10]), the system considered may have a non-square control gain matrix, and there may be faulty actuators that violate the controllability assumption imposed in these literature. The core technique of this paper is to decompose the control gain matrix and construct several auxiliary (not necessarily known) matrices to reconstruct the controllability of the system in the control design. The considered problem is well addressed by developing a framework for compatibility of this technique with PPC and backstepping. Notice that the proposed approach is different classical PPC since it does not rely on a closed-loop initialisation set and does not require control redesign and reanalysis for different performance requirements. Preliminary result of controllability relaxation for square nonlinear systems is presented in [16], where the classical PPC is considered. In this paper, we further relax the controllability condition for non-square systems subject to actuator faults, while eliminating the restrictions of the classical PPC so that prescribed performance tracking control with a unified framework can be achieved.
The main features and contributions of this work include:
  • The controllability conditions of MIMO systems with actuator faults have been relaxed where the control gain matrices are allowed to be non-square and the control design requires less a priori knowledge about such matrices.
  • By employing practical matrix decomposition and several auxiliary matrices, we propose a design framework applicable to both square and non-square systems, which paves the way for solving the global prescribed performance tracking control problem for MIMO nonlinear systems.
  • It is shown that the proposed controller exhibits better robustness in the sense of actuator failures and uncertain control gains compared to those in the existing literature (see, for instance, [12, 15]). In other words, the developed controller is applicable to the cases with or without actuator failures in the aforementioned literature and vice versa.
@@ II Problem Statement
$\dot{\bm{x}}_{i}=$ $\bm{f}_{i}(\bm{X}_{i})+\bm{g}_{i}(\bm{X}_{i},t)\bm{x}_{i+1}+{\bm{d}_{i}}(\bm{X}_{i}),i=1,\ldots,N-1,$
$\dot{\bm{x}}_{N}=$ $\bm{f}_{N}(\bm{X}_{N})+\bm{g}_{N}(\bm{X}_{N},t)\bm{u}_{a}+{\bm{d}_{N}}(\bm{X}_{N}),$
$\bm{y}=$ $\bm{x}_{1},$ (1)
where $\bm{x}_{i}=[x_{i1},\ldots,x_{in}]^{T}\in\mathbb{R}^{n}$ ($i=1,\ldots,N$) is the $i$th state vector; $\bm{X}_{i}=[\bm{x}_{1}^{T},\ldots,\bm{x}_{i}^{T}]^{T}\in\mathbb{R}^{n\times i}$; ${\bm{y}}=[y_{1},\ldots,y_{n}]^{T}\in\mathbb{R}^{n}$ is the output vector; ${\bm{f}_{i}}(\bm{X}_{i}):\mathbb{R}^{in}\rightarrow\mathbb{R}^{n}$ is some nonlinear uncertainty; ${\bm{d}_{i}}(\bm{X}_{i}):\mathbb{R}^{in}\rightarrow\mathbb{R}^{n}$ denotes external disturbance; $\bm{u}_{a}=[u_{a1},\ldots,u_{am}]^{T}\in\mathbb{R}^{m}$ is the input vector (the output of actuators); ${\bm{g}_{i}}(\bm{X}_{i}):\mathbb{R}^{in}\rightarrow\mathbb{R}^{n\times m}$, $n\leq m$ is an unknown control gain matrix.
@@
    $\mathcal{H}(-\underline{\delta}_{j}\varphi_{j}(t))<e_{j}(t)<\mathcal{H}(\bar{\delta}_{j}\varphi_{j}(t)),\ j=1,2,\ldots,m,$ (3)
@@
$\bm{g}_{i}=\bm{A}\bm{b}_{i},n<m,\ i=1,\ldots,N$ (4)
where $\bm{A}=[\bm{I}_{n},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{(m-n)}]\in\mathbb{R}^{n\times m}$ is a known constant matrix with $\bm{\Lambda}_{j}=[\iota_{j1},\ldots,\iota_{jn}]^{T}\in\mathbb{R}^{n}$ ($j=1,\ldots,m-n$) being a nonzero vector and $\iota_{jp}\geq 0$ ($p=1,\ldots,n$), and $\bm{{b}}_{i}\in\mathbb{R}^{m\times m}$ is an extension matrix of $\bm{g}_{i}$ which is completely unknown. Without loss of generality, we assume ${\bm{g}_{k}}\in\mathbb{R}^{n\times n}$ is square ($k=1,\ldots,N-1$) and ${\bm{g}_{N}}\in\mathbb{R}^{n\times m}$ with $n<m$ is non-square.
@@
Assumption 3 ensures the controllability of system (1), which is milder than the current state-of-the-art [6, 13, 10, 20, 16]. In [13] and [20], similar conditions are considered, yet $\bm{P}_{i}$ is required to be known and diagonal form, respectively, and the condition in [16] is only available for square systems. Thus, the controllability conditions in [6, 13, 16] can be viewed as some special cases of Assumption 3 since $\bm{P}_{i}$ is allowed to be unknown and can handle non-square nonlinear systems in our case. Moreover, it is worth noting that the auxiliary matrix $\bm{P}_{i}$ is only required for stability analysis, as opposed to needing to be present in the control design as in [13], which thus benefits the implementation of our method.
@@
In contrast to the decomposition of $\bm{g}_{N}$ in [2] and [9], no information about $\bm{g}_{N}$ except the structural features is required in Assumption 2, and the decomposed known matrix $\bm{A}$ is skilfully constructed to be uniform and simple for all systems, and does not prevent the attainment of the controllability condition. In [2] and [9], the controllability condition is established by decomposing the control gain matrix to obtain a system information-based matrix $\bm{L}\in\mathbb{R}^{n\times m}$ with full-row rank, however, such a decomposition is complex and hard to implement for some systems, moreover, such a controllability condition may be invalidated due to the presence of actuator faults (see Example 2 for detail). It is worth emphasizing that by Assumptions 2 and 3, the controllability of the system is ensured by matching the auxiliary matrix $\bm{P}_{i}$ analytically, rather than by solving the matrix $\bm{L}$ in a complex offline manner, as seen in our later development.
In what follows, we present two examples to shed light on the merits of Assumption 2 and Assumption 3.
Example 1: (square case) Consider
@@
It is trivial to verify that neither $\bm{g}$ nor $\bm{g}+\bm{g}^{T}$ are SPD. One can choose an auxiliary matrix $\bm{P}$ as follows
$\bm{P}=\begin{bmatrix}3&0\\ 0&1\end{bmatrix},\ \textrm{then}\ \bm{P}\bm{g}+\bm{g}^{T}\bm{P}=\begin{bmatrix}12&8\\ 8&6\end{bmatrix},$
which shows that Assumption 3 is satisfied and the controllability of the system is guaranteed.
Example 2: (non-square case) Consider a spacecraft equipped with four reaction wheels (RWs) [19], and the control gain matrix is defined as $\bm{g}=\bm{J}^{-1}\bm{D}\in\mathbb{R}^{3\times 4}$, where the inertia matrix $\bm{J}\in\mathbb{R}^{3\times 3}$ and configuration matrix $\bm{D}\in\mathbb{R}^{3\times 4}$ are given by
$\bm{J}=\begin{bmatrix}20&1.2&0.9\\ 1.2&5&1.4\\ 0.9&1.4&5\end{bmatrix}\textrm{kg}\cdot\textrm{m}^{2},\ \bm{D}=\begin{bmatrix}1&0&0&{\sqrt{3}}\\ 0&1&0&{\sqrt{3}}\\ 0&0&1&{\sqrt{3}}\end{bmatrix}.$
Assume that the actuation effectiveness matrix is set as $\bm{\rho}(t)=\textmd{diag}\{0.01,0.1+0.01\sin(t),0.01,0.1+0.01\cos(t)\}$. If we choose $\bm{L}=\bm{D}$ as in [21] and [19], by a simple calculation, it can be verified that $\bm{G}^{*}={\bm{{g}}\bm{\rho}\bm{D}^{T}+\bm{D}\bm{\rho}\bm{{g}}^{T}}$ is not positive definite, but we can still make Assumption 3 hold by choosing
$\bm{A}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&1\end{bmatrix}\ \textrm{and}\ \bm{P}=\begin{bmatrix}0.7+0.1\sin(t)&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}.$
@@
$\|\bm{f}_{i}(\bm{X}_{i})+{\bm{d}_{i}}(\bm{X}_{i})\|\leq a_{fi}\phi_{fi}(\bm{X}_{i}),$ (5)
$\|\bm{P}_{1}(t)\|\leq a_{11},\ \big{\|}\frac{\partial\bm{P}_{1}(t)}{\partial t}\big{\|}\leq a_{12},$ (6)
@@
$\big{\|}\frac{\partial\bm{P}_{k}(\bm{X}_{k-1},t)}{\partial t}\big{\|}\leq a_{k2}\phi_{k2}(\bm{X}_{k}),$ (7)
@@ III-A Non-monotonic Performance Function
To achieve the prescribed tracking performance, we first introduce the rate function $\beta(t)$ satisfying the following conditions: 1) $\beta(0)=1$; 2) $\beta(t)\in[0,1),\forall t>0$; and 3) $\beta^{(i)},i=0,1,...,N$, is known, bounded, and piecewise continuous.
@@
$\varphi(t)=(\varphi_{0}-\varphi_{f})\beta(t)+\varphi_{f},$ (8)
where $\varphi_{f}$ and $\varphi_{0}$ are some design parameters and satisfy $0<\varphi_{f}<\varphi_{0}\leq 1$. It is readily shown that $\varphi(t)$ exhibits the following properties: 1) $\varphi(0)=\varphi_{0}$ and $\varphi:(0,\infty)\rightarrow[\varphi_{f},\varphi_{0})$ for $t>0$; and 2) $\varphi^{(i)},i=0,1,...,n,$ is bounded, known, and piecewise continuous over the interval $[0,\infty)$.
Definition 1: A continuous function $\mathcal{H}:(-b_{0},b_{0})\rightarrow[-\infty,\infty)$ is called a intermediate function if 1) $\mathcal{H}$ is continuously differentiable; 2) $\mathcal{H}(\varsigma)=\mathcal{H}(-\varsigma)$, $\mathcal{H}(0)=0$, and $\lim_{\varsigma\rightarrow\pm b_{0}}\mathcal{H}(\varsigma)=\pm\infty$ ; and 3) $0<\upsilon\leq\dot{\mathcal{H}}$ and $\lim_{\varsigma\rightarrow\pm b_{0}}\dot{\mathcal{H}}(\varsigma)=\infty$, where $\upsilon$ is a constant.
@@
According to the definitions of $\varphi$ and $\mathcal{H}$, we call $\mathcal{H}(\varphi)$ the performance function, and it is trivial to prove that $\mathcal{H}$ is a non-monotonic function as $\varphi$ is a non-monotonic function. To facilitate the presentation, this paper adopts the non-monotonic performance function as in [15]
$\mathcal{H}(\varphi)=\frac{l\varphi}{\sqrt{1-\varphi^{2}}}\ \textrm{with}\ \beta(t)=\exp(-\gamma t)\cos^{2}(t),$ (9)
where $l$ and $\gamma$ are some positive constants. Note that in [8, 12], the scaling function $\varphi(t)$ is monotonically decreasing in time, however, as mentioned in [22] and [15], utilizing a non-monotonic function that can widen the performance boundary over certain time intervals may be beneficial in some practical scenarios.

### v2

@@ Abstract
In this paper, we investigate the problem of unified prescribed performance tracking for a class of non-square strict-feedback nonlinear systems under relaxed controllability conditions. By using a skillful matrix decomposition and introducing some feasible auxiliary matrices, a more generalized controllability condition than the current state of the art is constructed, which can be applied to both square and non-square nonlinear systems subject to actuator faults and unknown yet time-varying control gain. Incorporating the relaxed controllability conditions and the uniform performance specifications into the backstepping design procedure, a prescribed performance fault-tolerant controller is developed that can achieve different performance demands without modifying the controller structure, which is more flexible and practical. In addition, the destruction of the system stability by unknown controllability auxiliary matrices and unknown nonlinearities is circumvented by embedding the available core information of the state-dependent uncertainties into the design procedure. Both theoretical analysis and numerical simulation demonstrate the effectiveness and benefits of the proposed method.
@@
Non-square nonlinear systems, controllability condition, prescribed performance, actuator fault.
@@ I Introduction
Inspired by the above discussion, in this paper, we focus on solving the problem of prescribed performance control (PPC) of strcit-feedback MIMO systems under a more relaxed controllability condition. Unlike the existing literatures on PPC for MIMO systems (e.g., [8, 10, 16]), the system considered may have a non-square control gain matrix, and there may be faulty actuators that violate the controllability assumption imposed in these literature. The core technique of this paper is to decompose the control gain matrix and construct several auxiliary (not necessarily known) matrices to reconstruct the controllability of the system in the control design. The considered problem is well addressed by developing a framework for compatibility of this technique with PPC and backstepping. It is worth noting that the proposed approach differs from classical PPC methods in that it does not rely on a closed-loop initialization set, nor does it require control redesign or reanalysis when performance requirements change. Preliminary result of controllability relaxation for square nonlinear systems is presented in [17], where the classical PPC is considered. In this paper, we further relax the controllability condition for non-square systems subject to actuator faults, while eliminating the restrictions of the classical PPC so that prescribed performance tracking control with a unified framework can be achieved. The main features and contributions of this work include:
  1. (1)
    We construct a more general controllability condition than those in existing approaches [17, 18, 2], enabling the proposed method to accommodate a broader class of system models, including both square and non-square cases with unknown, time-varying control gain matrices, even in the presence of intermittent actuator faults.
  2. (2)
    The proposed PPC method supports more general and practical performance specifications without conflicting with the relaxed controllability conditions. Unlike existing methods [8, 12, 19, 20], which typically handle only symmetric performance functions, our approach enables both symmetric and asymmetric behaviors through simple tuning of design parameters, without modifying the control structure or reanalyzing stability.
  3. (3)
    In contrast to previous works [12, 15], we embed the auxiliary matrix, rather than the actuation effectiveness matrix, into the Lyapunov function. By exploiting the state-dependent structure in the backstepping design, this approach overcomes stability challenges arising from the generalized controllability conditions while ensuring boundedness of all closed-loop signals.
@@ II Problem Statement
$\dot{\bm{x}}_{k}=\$$\bm{f}_{k}(\bm{X}_{k})+\bm{g}_{k}(\bm{X}_{k},t)\bm{x}_{k+1}+{\bm{d}_{k}}(\bm{X}_{k}),$
$\dot{\bm{x}}_{N}=\$$\bm{f}_{N}(\bm{X}_{N})+\bm{g}_{N}(\bm{X}_{N},t)\bm{u}_{a}+{\bm{d}_{N}}(\bm{X}_{N})$
$\bm{y}=\$$\bm{x}_{1},\ \ k=1,\ldots,N-1$ (1)
where $\bm{x}_{i}=[x_{i1},\ldots,x_{in}]^{T}\in\mathbb{R}^{n}$ ($i=1,\ldots,N$) is the $i$th state vector; $\bm{X}_{i}=[\bm{x}_{1}^{T},\ldots,\bm{x}_{i}^{T}]^{T}\in\mathbb{R}^{n\times i}$; ${\bm{y}}=[y_{1},\ldots,y_{n}]^{T}\in\mathbb{R}^{n}$ is the output vector; ${\bm{f}_{i}}:\mathbb{R}^{in}\rightarrow\mathbb{R}^{n}$ is some nonlinear uncertainty; ${\bm{d}_{i}}:\mathbb{R}^{in}\rightarrow\mathbb{R}^{n}$ denotes external disturbance; $\bm{u}_{a}=[u_{a1},\ldots,u_{am}]^{T}\in\mathbb{R}^{m}$ is the input vector (the output of actuators); ${\bm{g}_{i}}:\mathbb{R}^{in}\rightarrow\mathbb{R}^{n\times m}$, $n\leq m$ is an unknown control gain matrix.
@@
    $\mathcal{H}(-\underline{\delta}_{j}\varphi_{j}(t))<e_{j}(t)<\mathcal{H}(\bar{\delta}_{j}\varphi_{j}(t)),\ j=1,2,\ldots,m$ (3)
@@
$\bm{g}_{i}(\bm{X}_{i},t)=\bm{A}\bm{b}_{i}(\bm{X}_{i},t),n<m,\ i=1,\ldots,N$ (4)
where $\bm{A}=[\bm{I}_{n},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{(m-n)}]\in\mathbb{R}^{n\times m}$ is a known actuator allocation matrix, with $\bm{\Lambda}_{j}=[\iota_{j1},\ldots,\iota_{jn}]^{T}\in\mathbb{R}^{n}$ being a nonzero vector satisfying $\iota_{jp}\geq 0$ ($j=1,\ldots,m-n$, $p=1,\ldots,n$), and $\bm{{b}}_{i}\in\mathbb{R}^{m\times m}$ is an extension matrix of $\bm{g}_{i}$. Without loss of generality, assume ${\bm{g}_{k}}\in\mathbb{R}^{n\times n}$ ($k=1,\ldots,N-1$) and ${\bm{g}_{N}}\in\mathbb{R}^{n\times m}$ with $n<m$.
@@
Assumption 3 ensures the controllability of system (1), which is milder than the current state-of-the-art [6, 13, 10, 24, 25, 17]. In [13] and [25], similar conditions are considered, yet $\bm{P}_{i}$ is required to be known and diagonal form, respectively, and the condition in [17] is only available for square systems. Thus, the controllability conditions in [6, 13, 17] can be viewed as some special cases of Assumption 3 since $\bm{P}_{i}$ is allowed to be unknown and can handle non-square nonlinear systems in our case. Moreover, it is worth noting that the auxiliary matrix $\bm{P}_{i}$ is only required for stability analysis, as opposed to needing to be present in the control design as in [13], which thus benefits the implementation of our method.
@@
The decomposition in Assumption 2 represents a simplified input-output mapping, where $\bm{I}_{n}$ denotes the primary input contribution, and $\bm{\Lambda}_{j}$ captures the effect of additional inputs, such as redundant actuators or extra degrees of freedom. The specific form of $\bm{\Lambda}_{j}$ in $\bm{A}$ depends on the system’s redundant dynamic allocation. For example, $\bm{\Lambda}_{j}=[1/\sqrt{3},\,1/\sqrt{3},\,1/\sqrt{3}]^{T}$ implies that the redundant input contributes equally to all primary channels, whereas $\bm{\Lambda}_{j}=[0,\,0,\,1]^{T}$ indicates that it only affects one of them. Such decomposition strategies are commonly employed in redundant control systems, including robotics and spacecraft [26]. It is worth emphasizing that the adopted decomposing strategy is only valid for over-actuated systems with $n<m$. Under this condition, as long as $\bm{A}$ is full row rank, the decomposition can represent any $\bm{g}\in\mathbb{R}^{n\times m}$ without loss of generality, thereby avoiding modeling errors or structural limitations. In contrast, for the under-actuated case $n>m$, the column space of $\bm{A}$ is restricted to at most an $m$-dimensional subspace of $\mathbb{R}^{n}$, which limits the admissible range of $\bm{g}$ and prevents the coverage of all possible configurations. This structural limitation undermines both controllability analysis and controller design generality. Therefore, to avoid such degeneracy and potential singularities in the modeling framework, this work exclusively focuses on non-square MIMO systems with redundant inputs ($n<m$), under which the proposed control strategy and the associated extended controllability conditions are developed.
Remark 3
Unlike [2] and [9], where $\bm{g}_{N}$ is explicitly decomposed, Assumption 2 requires no additional information about $\bm{g}_{N}$ beyond structural characteristics. That is, $\bm{A}$ only reflects predefined dynamic allocation and does not ensure system controllability. In [2] and [9], the controllability condition is established by decomposing the control gain matrix to obtain a system information-based matrix $\bm{L}\in\mathbb{R}^{n\times m}$ with full row rank, however, such a decomposition is complex and hard to implement for some systems, moreover, such a controllability condition may be invalidated due to the presence of actuator faults (see Example 2 for detail). It is worth emphasizing that by Assumptions 2 and 3, the controllability of the system is ensured by matching the auxiliary matrix $\bm{P}_{i}$ analytically, rather than by solving the matrix $\bm{L}$ in a complex offline manner, as seen in our later development.
To validate the effectiveness of Assumption2 and Assumption3, two examples are presented to show how they facilitate the relaxation of standard controllability assumptions and improve the applicability of the control design.
Example 1
(Square case) Consider
@@
It is easy to confirm that neither $\bm{g}$ nor $\bm{g}+\bm{g}^{T}$ are SPD, which implies that the conventional controllability assumptions employed in [6, 10] do not hold in this case. However, by selecting an appropriate auxiliary matrix
$\bm{P}=\begin{bmatrix}3&0\\ 0&1\end{bmatrix},\ \textrm{then}\ \bm{P}\bm{g}+\bm{g}^{T}\bm{P}=\begin{bmatrix}12&8\\ 8&6\end{bmatrix}$
which is symmetric positive definite. This shows that Assumption 3 holds, thus guaranteeing the controllability of the system under the relaxed condition.
Example 2
(Non-square case) Consider a spacecraft equipped with four reaction wheels (RWs) [23], and the control gain matrix is defined as $\bm{g}=\bm{J}^{-1}\bm{D}\in\mathbb{R}^{3\times 4}$, where the inertia matrix $\bm{J}\in\mathbb{R}^{3\times 3}$ and configuration matrix $\bm{D}\in\mathbb{R}^{3\times 4}$ are given by
$\bm{J}=\begin{bmatrix}20&1.2&0.9\\ 1.2&5&1.4\\ 0.9&1.4&5\end{bmatrix}\textrm{kg}\cdot\textrm{m}^{2},\ \bm{D}=\begin{bmatrix}1&0&0&{1/\sqrt{3}}\\ 0&1&0&{1/\sqrt{3}}\\ 0&0&1&{1/\sqrt{3}}\end{bmatrix}$.
Assume that the actuation effectiveness matrix is set as $\bm{\rho}(t)=\textmd{diag}\{0.1-0.08\sin(t),0.2-0.17\sin(t),0.2-0.15\sin(t),0.2+0.18\sin(t)\}$. If we choose $\bm{A}=\bm{D}$ as in [26] and [23], by a simple calculation, it can be verified that $\bm{G}^{*}={\bm{{g}}\bm{\rho}\bm{D}^{T}+\bm{D}\bm{\rho}\bm{{g}}^{T}}$ is not positive definite, but we can still make Assumption 3 hold by choosing
$\bm{P}=\begin{bmatrix}0.7+0.1\sin(t)&0&0\\ 0&0.1&0\\ 0&0&0.6+0.1\cos(t)\end{bmatrix}$.
Consequently, the system maintains controllability under actuator degradation and varying input effectiveness, demonstrating the robustness and flexibility of the proposed framework for non-square, over-actuated MIMO systems such as spacecraft with multiple reaction wheels.
@@
$\|\bm{f}_{i}(\bm{X}_{i})+{\bm{d}_{i}}(\bm{X}_{i})\|\leq a_{fi}\phi_{fi}(\bm{X}_{i})$ (5)
$\|\bm{P}_{1}(t)\|\leq a_{11},\ \big{\|}\frac{\partial\bm{P}_{1}(t)}{\partial t}\big{\|}\leq a_{12}$ (6)
@@
$\big{\|}\frac{\partial\bm{P}_{k}(\bm{X}_{k-1},t)}{\partial t}\big{\|}\leq a_{k2}\phi_{k2}(\bm{X}_{k})$ (7)
@@ III-A Non-monotonic Performance Function
To achieve the prescribed tracking performance, we first introduce the rate function $\beta(t)$ satisfying the following conditions: 1) $\beta(0)=1$; 2) for all $t>0$, $\beta(t)$ remains within the interval $[0,1)$; and 3) all derivatives $\beta^{(i)}$ for $i=0,1,\dots,N$ are known, uniformly bounded, and piecewise continuous over $0,\infty)$.
@@
$\varphi(t)=(\varphi_{0}-\varphi_{f})\beta(t)+\varphi_{f}$ (8)
where $\varphi_{f}$ and $\varphi_{0}$ are some design parameters and satisfy $0<\varphi_{f}<\varphi_{0}\leq 1$. It can be verified that $\varphi(t)$ satisfies the following characteristics: 1) $\varphi(0)=\varphi_{0}$ and $\varphi:(0,\infty)\rightarrow[\varphi_{f},\varphi_{0})$ for $t>0$; and 2) all derivatives $\varphi^{(i)}$ for $i=0,1,\dots,n$ are known, uniformly bounded, and piecewise continuous on $0,\infty)$.
Definition 1
A continuous function $\mathcal{H}:(-b_{0},b_{0})\rightarrow[-\infty,\infty)$ is called a intermediate function if 1) $\mathcal{H}$ is continuously differentiable; 2) $\mathcal{H}(\varsigma)=-\mathcal{H}(-\varsigma)$, $\mathcal{H}(0)=0$, and $\lim_{\varsigma\rightarrow\pm b_{0}}\mathcal{H}(\varsigma)=\pm\infty$ ; and 3) $0<\upsilon\leq\dot{\mathcal{H}}$ and $\lim_{\varsigma\rightarrow\pm b_{0}}\dot{\mathcal{H}}(\varsigma)=\infty$, where $\upsilon$ is a constant.
@@
According to the definitions of $\varphi$ and $\mathcal{H}$, we call $\mathcal{H}(\varphi)$ the performance function, and it is trivial to prove that $\mathcal{H}$ is a non-monotonic function as $\varphi$ is a non-monotonic function. To facilitate the presentation, this paper adopts the non-monotonic performance function as
$\mathcal{H}(\varphi)=\frac{l\varphi}{\sqrt{1-\varphi^{2}}}\ \textrm{with}\ \beta(t)=\exp(-\gamma t)\cos^{2}(t)$ (9)
where $l$ and $\gamma$ are some positive constants. As noted in [8, 12, 27], the scaling function $\varphi(t)$ is commonly designed to decrease monotonically over time. However, [28] and [15] have pointed out that employing a non-monotonic $\varphi(t)$, capable of temporarily relaxing the performance constraint, may offer practical benefits in specific scenarios.